ADNAN MUHAMMED

Grade 12,

Prove that , (sinX) 3 × sin3X + (cosX) 3 × cos3X =(cos2X) 3

Prove that ,
(sinX)3 × sin3X + (cosX)3 × cos3X =(cos2X)3

Grade:12th pass

1 Answers

Susmita
5 years ago
Use the formula
sin3x=(3sinx-sin3x)/4
and cos3x=(3cosx+cos3x)/4
also cos3(2x)=(3cos2x+cos6x)/4
Put the first two formula in the given expression.you will get
(3sinxsin3x-sin23x+3cosxcos3x+cos23x)/4
Put 3sinxsin3x+3cosxcos3x=cos(3x-x)=cos2x
And cos23x-sin23x=cos(2.3x)=cos6x.
and see you will get the ans.